FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF N-METHYLPROPIONAMIDE

YOSHIYUKI KAWASHIMA, Department of Applied Chemistry, Kanagawa Institute of Technology, Atsugi, Kanagawa 243-0292, JAPAN; RICHARD D. SUENRAM, National Institute of Standards and Technology, Gaithersburg, MD 20899; and \underline{EIZI HIROTA}, The Graduate University for Advanced Studies, Hayama, Kanagawa 240-0193, JAPAN.

In order to clarify the dynamical behavior of the peptide bond, we have undertaken a systematic study of ′ peptide molecules ′, which consist of (a) peptide bond(s) with internal-rotation groups at the both ends of the bond(s). In the present investigation we focused attention to the molecule (NMPA) shown in the title, which has an ethyl group (CH_{3} of which is referred to as C-CH_{3}) at the carbonyl side and a methyl group (called N-CH_{3}) at the amide side, and aimed at unveiling how the two CH_{3} groups interact with each other through the peptide bond. We have derived a rotational Hamiltonian including the two CH_{3} internal rotations, and have treated the C-CH_{3} internal rotation by a conventional PAM, while applying a more sophisticated approach to the N-CH_{3} internal rotation. NMPA may be regarded to belong to group G_{18}, even if its skeleton executes large-amplitude ′ out-of-plane ′ motions. The group consists of 6 species: A_{1}, A_{2}, E_{1}, E_{2}, E_{3}, and E_{4}. We have observed and analyzed A_{1} (or A_{2}) and E_{2} spectra, but have not detected any lines due to the first excited state of the CH_{3}CH_{2}-CO torsion of A_{2} symmetry, indicating that the internal-rotation splitting is quite large. The potential barrier to the C-CH_{3} internal rotation was determined to be 799 cm^{−1}, which may be compared with that of N-CH_{3} of about 81 cm^{−1}. The coupling between the two CH_{3}’s is being analyzed by observing E_{1}, E_{3}, and E_{4} spectra.